OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x

OpenStudy (anonymous):

sin2x = 2sinxcosx cos2x = cos^2x-sin^2x = (cosx - sinx)(cosx + sinx)

OpenStudy (anonymous):

Ok... I'm confused as to what comes next? 2sinxcosx-(cosx-sinx)(cosx+sinx) ?

OpenStudy (anonymous):

@findme

OpenStudy (anonymous):

doesn't that answer your question (rewrite the expression with only sinx and cosx)?

OpenStudy (anonymous):

My answer choices are 2 sin2x - 2 sin x cos x + 1 2 sin x 2 sin2x + 2 sin x cos x - 1 2 sin2x - 2 sin x cos x - 1

OpenStudy (anonymous):

2 sin^2x - 2 sin x cos x + 1 2 sin x 2 sin^2x + 2 sin x cos x - 1 2 sin^2x - 2 sin x cos x - 1

OpenStudy (anonymous):

cos2x = 1 - 2sin^2x how about now? can you figure it out?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

sin2x = 2sinxcosx cos2x = 1 - 2sin^2x so, substituting the above for the given expression: sin2x - cos2x = 2sinxcosx - (1 - 2sin^2x) = 2sin^2x + 2sinxcosx - 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np